OpenStudy (anonymous):

The three ways of solving Quadratic equations.?? some dont seem to be able to factorise. You cant always find the products to add to get the right answers, so is this the time you use the the standard quadratic equation ? hElP :)

hartnn (hartnn):

which ways you already know ?

OpenStudy (anonymous):

usually, I find the products of the last term that will add to satisfy the other x term, works nearly all the time, gets a tiny bit , well one more step if the coefficient of x is not 1 but no problems. Some just dont get enough products to add to make anywhere near the correct factorization, so is that where the quadratic formula has to be used ?

hartnn (hartnn):

use of Quadratic formula : Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\) Factorizing method : for factorizing the standard quadratic equation,\( ax^2+bx+c=0\), we need 2 numbers whose sum is 'b' and product = ac So that we can split the middle term Completing the square method : .....

hartnn (hartnn):

Quadratic Formula works EVERYTIME .

OpenStudy (anonymous):

yep im with you

hartnn (hartnn):

even when the quadratic expression can be factored, Quadratic formula and Completing the square method will work every time. Factor method will not work ONLY WHEN b^2-4ac is NOT a perfect square

hartnn (hartnn):

you know Completing the square method or should i explain ?

OpenStudy (anonymous):

thanks, i got the square method :) I usually find it quick but i was just unsure if i couldnt do it that way was the quad formula the only way to go.... thanks. that was great help !! :)

OpenStudy (anonymous):

oh, Ill do some more completing the square method. than i should be ok, thx.

hartnn (hartnn):

good :) welcome ^_^

OpenStudy (anonymous):

can you give one example of completing the square ?

OpenStudy (anonymous):

like a hrd hard one :)

hartnn (hartnn):

yup, let me try...

hartnn (hartnn):

\(2x^2+5x-9\)

hartnn (hartnn):

lol, couldn't think of harder :P

OpenStudy (anonymous):

yep factors of 18 that add to ummmm 5 ..no..umm

hartnn (hartnn):

you trying factor method ? you wanted completing the square method example, right ?

OpenStudy (anonymous):

oh..yes...whoops

OpenStudy (anonymous):

Ill get it done... just covered so much in the last week. thans, ill be alright.

hartnn (hartnn):

cool :)